hello,
my name is lauren and i am taking a algebra function and models class at suny westchter community college. I have and adjunct professor whos strong point is wall street, and here lies my problem. she has been telling the class that the equation to find a slope of a line is y1-y2 over x1-x2 and i know its y2-y1 over x2-x1, but she says it doesnt matter you ge the same answer either way but i keep getting diffrenet answers. Is she right and I am doing my math wrong or am i right that the formula she is giving is wrong? please help i need this class to graduate.

Answer by Alan3354(69443)

she has been telling the class that the equation to find a slope of a line is y1-y2 over x1-x2 and i know its y2-y1 over x2-x1, but she says it doesnt matter you ge the same answer either way but i keep getting diffrenet answers. Is she right and I am doing my math wrong or am i right that the formula she is giving is wrong? please help i need this class to graduate.
------------
She's not wrong.
It's the difference in y over the diff in x, as long as you use either of the points as (x1,y1)
--------------
(x1 - x2)/(y1 - y2) = -(x2 - x1)/-(y2 - y1) = (x2 - x1)/(y2 - y1)
Same thing.